# LFTC-25-11 / 105

### v1

@@ Acknowledgements
The following conventions are used for the LF components of the co-ordinate 4-vector $x^{\mu}$ in the present work: $x^{+}=\frac{x^{0}+x^{3}}{\sqrt{2}}$ (time coordinate), $x^{-}=\frac{x^{0}-x^{3}}{\sqrt{2}}$ (longitudinal space coordinate) and ${\bf{x}}^{\perp}=(x^{1},x^{2})$ (transverse space coordinates).
The conjugate LF components of the momentum 4-vector $p^{\mu}$ are $p^{-}=\frac{p^{0}-p^{3}}{\sqrt{2}}$ (energy), $p^{+}=\frac{p^{0}+p^{3}}{\sqrt{2}}$ (longitudinal momentum) and ${\bf{p}}^{\perp}=(p^{1},p^{2})$ (transverse momentum).
@@
The $\gamma$-matrices are given by the following representation
@@
The $\gamma$-matrices satisfy the relations

### v2

@@ Acknowledgments
The following conventions are used for the LF components of the coordinate four-vector $x^{\mu}$ in the present work: $x^{+}=\frac{x^{0}+x^{3}}{\sqrt{2}}$ (time coordinate), $x^{-}=\frac{x^{0}-x^{3}}{\sqrt{2}}$ (longitudinal space coordinate) and ${\bf{x}}^{\perp}=(x^{1},x^{2})$ (transverse space coordinates).
The conjugate LF components of the momentum four-vector $p^{\mu}$ are $p^{-}=\frac{p^{0}-p^{3}}{\sqrt{2}}$ (energy), $p^{+}=\frac{p^{0}+p^{3}}{\sqrt{2}}$ (longitudinal momentum), and ${\bf{p}}^{\perp}=(p^{1},p^{2})$ (transverse momentum).
@@
The $\gamma$ matrices are given by the following representation
@@
The $\gamma$ matrices satisfy the relations